# Alberta Wells Dataset: Pinpointing Oil and Gas Wells from Satellite Imagery

### v2

@@ Abstract
Millions of abandoned oil and gas wells are scattered across the world, leaching methane into the atmosphere and toxic compounds into the groundwater. Many of these locations are unknown, preventing the wells from being plugged and their polluting effects averted. Remote sensing is a relatively unexplored tool for pinpointing abandoned wells at scale. We introduce the first large-scale benchmark dataset [^1] Dataset available at: https://zenodo.org/records/13743323 for this problem, leveraging medium-resolution multi-spectral satellite imagery from Planet Labs. Our curated dataset comprises over 213,000 wells (abandoned, suspended, and active) from Alberta, a region with especially high well density, sourced from the Alberta Energy Regulator and verified by domain experts. We evaluate baseline algorithms [^2] Code will be available at: https://github.com/RolnickLab/Alberta_Wells_Dataset for well detection and segmentation, showing the promise of computer vision approaches but also significant room for improvement.
@@ 1 Introduction
Across the world, there are millions of abandoned oil and gas wells left to degrade by the companies or individuals that built them. No longer producing usable fossil fuels, these wells nonetheless have a significant impact on the environment, with many of them leaking significant quantities of methane, a powerful greenhouse gas, into the atmosphere. In aggregate, these emissions represent the equivalent of millions of tons of carbon dioxide per year [1]. In Canada, an estimated 370,000 abandoned wells produce methane equivalent to half a million metric tons of CO2 annually [1, 2], while in the U.S., there are an estimated 4 million abandoned wells [1], releasing over five million metric tons of CO2 equivalent emissions per year.
Despite the small percentage, the high global warming potential of methane underscores the need for improved monitoring and management to reduce environmental impact. Abandoned wells also pose health and safety concerns, in particular by leaching toxic chemicals into the groundwater of surrounding communities [3]. It is possible to plug abandoned wells to mitigate the harms associated with them (with so-called “super-emitter” wells an especially high priority [4, 5]). However, a significant fraction of abandoned wells remain unknown. In Pennsylvania, as much as 90% of abandoned wells are estimated to be unrecorded [5]. In Canada, abandoned wells have been described as the most uncertain source of methane emissions nationally due to the poor quality of data surrounding them [1].
With the advent of large-scale remote sensing datasets and powerful machine learning tools to process them, it has become possible to label and monitor the built environment as never before [6]. Many such works have focused on opportunities to use remote sensing to accelerate climate action and environmental protection, and oil and gas infrastructure has increasingly been an object of scrutiny (see e.g. [7, 8]).
In this work, we present the first large-scale machine-learning benchmark dataset for pinpointing onshore oil and gas wells, encompassing abandoned, suspended, and active wells. Our main contributions are as follows:
  • We introduce the Alberta Wells Dataset, containing information on over 200k abandoned, suspended, and active onshore oil and gas wells with high-resolution satellite imagery.
@@
2 Previous Work
Hundreds of satellites continuously monitor the Earth’s surface, generating petabyte-scale remote sensing datasets [6]. With advancements in hardware, the quality of remote sensing images has significantly improved in terms of spatial and temporal resolution. High-quality remote sensing data are available through state-funded projects like Sentinel and Landsat, and more recently through private enterprises such as Planet Labs [9].
Increasingly, machine learning has been used to parse such raw data, including in a wide range of applications for tackling climate change [7]. Benchmark datasets in this area have included tasks in land use and land cover (LULC) estimation [10], crop classification [11, 12], species distribution modeling [13], and forest monitoring [14].Some datasets like SpaceNet 7 [15] include a few cases of study sites with oil wells, although the dataset was developed for multi-temporal urban monitoring.
Within this area of research, an increasing body of work has considered the problem of detecting artifacts associated with oil and gas operations. The detection of oil spills using a combination of remote sensing and machine learning has been widely explored [16, 17, 18]. Recently, the detection of oil and gas infrastructure has also been investigated [8, 19], with some studies focusing on the goal of estimating methane emissions [20, 21]. The dataset by [8] includes 7,066 aerial images, with 149 images of oil refineries. The METER-ML dataset [20] comprises 86,599 georeferenced images in the U.S. labeled for methane sources. The OGIM v1 dataset [21] includes 2.6 million point locations of major facilities. A dataset by [22] features 1,388 images of pipelines in the Arctic, while a dataset by [23] includes 3,266 images of heavy-polluting enterprises with 0.25 m resolution.
The problem of detection of oil and gas wells has also been proposed by a number of authors. Existing datasets, however, are quite small (500-10,000 samples) and typically are limited to a small region and contain only active wells, limiting their applicability in the context of identifying abandoned or suspended wells as summarized in Table 1. Most of these studies have primarily focused on basic machine learning algorithms for well detection due to the limited sample size.
@@ 2 Previous Work
NEPU-OWOD V1.0 [24] 1,192 432 0.41 Daqing City, China Google Earth
NEPU-OWOD V3.0 [25] 3,749 722 0.48 China & California
Oil Well Dataset [26] N/A 5,895 0.26 Daqing City, China
O&G Infrastructure [27] 630 930 0.15 - 1 Permian Basin, USA
Well Pad Dataset [28] 12,490 10,432 0.3-0.7
@@
NEPU-OWS V1.0 [29] N/A 1,200 10 Russia Sentinel-2
NEPU-OWSV2.0 [30] N/A 120 10/20/60 Austin, USA
@@ 3 Alberta Wells Dataset
We introduce the Alberta Wells Dataset for oil and gas well detection. The dataset is drawn from the province of Alberta, Canada, a region with the third-largest oil reserves in the world and a substantial number of oil and gas wells, many of which have been present for over a century. The entire province of Alberta (an area larger than the UK and Germany combined) encompasses a diverse range of geographical zones and is highly diverse for a landlocked region, including prairies, lakes, forests, and mountains. The dataset contains over 188,000 patches of satellite imagery acquired by Planet Labs [9], which we make publicly available, with around 94,000 containing wells covering more than 213,000 individual wells. Each patch is annotated with labels for both segmentation and bounding box localization. The annotations are based on data from the Alberta Energy Regulator [31], quality-controlled by domain experts.
To simulate real-world conditions, we ensure a varied density of wells per image, as highlighted in Figure LABEL:fig:figure1. We also include satellite imagery patches with no wells present from areas nearby to areas with wells, ensuring no overlap between the samples. This balanced dataset maintains an equal distribution of well and non-well images. Table 2 details the total sample count in each dataset split. It also mentions the count of oil wells in active, abandoned, and suspended states in each dataset split, emphasizing the dataset’s comprehensive nature.
3.1 Data Collection, Quality Control & Patch Creation
The Alberta Energy Regulator (AER) oversees the energy industry in the province, ensuring companies adhere to regulations as they develop oil and gas resources. AER publishes AER ST37 [31], a monthly list of all wells reported in Alberta, detailing their geographic location, mode of operation, license status, and type of product being extracted, among other attributes. This data provides a metadata (.txt) file and a .shp shape file, where each entry represents a unique geo-location point per site but often contains duplicates. However, this data cannot be used directly because the license status or mode of operation does not always correlate with the actual status of the well and often contains duplicates. Therefore, we work with domain experts to perform quality control as illustrated in Figure 3.
First, we remove duplicate entries from the well metadata, which often contain multiple instances of the same well identified by duplicate license numbers. We resolve these duplicates by retaining the most recent update. A similar approach is applied to the shapefile, where duplicates are resolved using the license date.
Afterward, we merge both datasets and filter the data, categorizing the wells as active, abandoned, or suspended based on criteria developed in consultation with domain experts, as shown in Table 3. We check for duplicate location coordinates in the dataset and resolve them by retaining the instance with the latest drill date.
Finally, we ensure all the well instances in the dataset are within Alberta’s boundaries. The raw metadata file has around 637,000 instances, while the surface hole geometry file has around 512,000 instances. After filtering and performing quality control on the datasets with domain experts, we have around 217,000 instances. We then calculate the geographical bounds covered by the well instances across the province and divide the region into non-overlapping square image patches, each covering an area of 1.1025 sq km (with sides of 1050m).
These images include various numbers of individual wells (see Fig. LABEL:fig:figure1), and we ensure that an approximately equal number of patches exist with and without wells. As a result of this process, some samples were excluded due to being located outside Alberta’s geographical boundaries, leading to a final total of approximately 213,000 well instances in the dataset patches.
Figure 3: Flowchart depicting our process for AER ST37 dataset cleaning and quality control.
@@ 3.2 Dataset Splitting
To create a well-distributed dataset that represents various geographical regions and offers a diverse dataset for evaluation and testing, we developed a splitting algorithm (see Algorithm 1). Our splitting approach focuses on balancing regions, not individual examples, ensuring that both the training and test sets reflect a diverse range of regions from Alberta’s varied landscape. This approach preserves diversity and simulates real-world conditions where imbalances are common.
This method involves forming small clusters $k_{1i}$ of nearby well patches based on their centroids as illustrated in Figure LABEL:fig:figuresplit (a). These small clusters are then grouped into larger, non-intersecting super-clusters $k_{2i}$, with each super-cluster representing a city or larger geographical area. The formation of super-clusters involves calculating a centroid for each $k_{1i}$ cluster based on the centroids of the well patches it contains as illustrated in Figure LABEL:fig:figuresplit (b). By clustering wells in this manner, we ensure that $k_{1i}$ clusters group wells from nearby localities together, while $k_{2i}$ clusters group wells from the same geographic region as illustrated in Figure LABEL:fig:figuresplit (c). Thus, each $k_{2i}$ cluster represents a geographic distribution, with each $k_{1i}$ cluster within it representing a sample of that distribution.
To ensure a diverse and well-distributed evaluation and testing of our machine learning model, we select the $k_{1i}$ clusters with the two fewest well instances from each $k_{2i}$ super-cluster for inclusion in the evaluation and test sets. This approach ensures a diverse representation of the dataset as observed in Figure LABEL:fig:figuresplit (d).
@@
Moreover, we maintain an equal distribution of well and non-well patches. In cases of imbalance in non-well images, we exclude such patches from the contributing $k_{1i}$ clusters as specified in Algorithm 1. For imbalances in well images, we sample non-well patches that are not part of any other clusters.
@@ 3.3 Satellite Imagery Acquisition & Label Creation
We used PlanetScope-4-Band imagery [9] featuring RGB and Near Infrared bands to represent satellite images of the region with a medium resolution of about 3 meters per pixel. PlanetScope, a product of Planet Labs, consists of approximately 130 satellites that can image the entire Earth’s land surface daily, collecting up to 200 million sq. km of data each day. We obtained Surface Reflectance imagery, which is offset-corrected, flat-field-corrected, ortho-rectified, visually processed, and radio-metrically corrected. These processes ensure consistency across varying atmospheric conditions and minimize uncertainty in spectral response over time and location, making the data ideal for various applications. We choose Planet Labs data over other alternatives since it is updated daily, making it possible to pick a consistent time for all the images, which is important for training dataset consistency. It also provides multispectral imagery (4-band: RGB+Near Infrared), and the Near Infrared band is a useful addition (as we later verify in Table 6) since certain features, like ground depressions indicating well sites, may be more detectable in this band. Lastly, while other alternatives may be limited in remote regions, Planet’s global satellite constellation ensures more consistent coverage. All the imagery we use is made publicly available in the dataset. To ensure the highest quality, we selected images with no cloud cover. The images were acquired by Planet satellites within a timeframe that aligns with the well-location data from AER. We obtained satellite images for each sample based on geographical coordinates, ensuring an intersection between the actual area of interest and the acquired imagery.
We frame the task of identifying wells as both an object detection and segmentation task. In a real-world setting, high performance in either detection or segmentation would be sufficient for detection of wells; therefore we test both framings since it is not a priori clear which may be best suited to remote sensing algorithms. For each image patch, as shown in Figure 7, we generated corresponding segmentation maps and object detection annotations for all known wells in the image based on the point labels provided in the AER data.
@@
For binary segmentation, we annotated each well site with a circle to match the teardrop shape typical for well sites. We standardized the diameter of a well site to a value of 90 meters (such sites typically range from 70 to 120 meters in diameter). We used the same scale to define bounding boxes in the object detection task, following the COCO [32] format for annotations. The overlap in ground truth bounding boxes for some of the wells in Figure 7 and Figure 14 reflects the clustering of multiple wells in densely developed oil and gas sites, where the spatial overlap of wells and infrastructure is common. (Note that this is a characteristic of the data, not a limitation of our quality control strategy.) Additionally, we created multi-class segmentation maps, where each class represents a different state of the well (active, suspended, or abandoned), and included this information in the object detection annotations. (We do not perform multi-class segmentation experiments here, but it is possible that future researchers may find this task useful.)
Identifying wells at 3m resolution seems like a very difficult task That said, it’s important to consider that ML models may actually have the potential to surpass human performance in this particular task, for several reasons. First, identifying wells at a 3m resolution from satellite imagery is extremely challenging and requires significant patience by a human. Additionally, there are few human experts in this field, as well identification typically relies on existing databases or on-the-ground field studies rather than detailed satellite image analysis. Finally, certain features (such as depressions in the ground representing well sites) may be easier to pick out in the infrared channel, which humans are not accustomed to using, as compared with the RGB channels.

### v3

@@ Abstract
Millions of abandoned oil and gas wells are scattered across the world, leaching methane into the atmosphere and toxic compounds into the groundwater. Many of these locations are unknown, preventing the wells from being plugged and their polluting effects averted. Remote sensing is a relatively unexplored tool for pinpointing abandoned wells at scale. We introduce the first large-scale Benchmark Dataset for this problem, leveraging high-resolution multi-spectral satellite imagery from Planet Labs. Our curated Dataset comprises over 213,000 wells (abandoned, suspended, and active) from Alberta, a region with especially high well density, sourced from the Alberta Energy Regulator and verified by domain experts. We evaluate baseline algorithms for well detection and segmentation, showing the promise of computer vision approaches but also significant room for improvement.
Remote Sensing, Satellite Imagery, Climate Change, AI for Good, ICML
@@ 1 Introduction
Across the world, there are millions of abandoned oil and gas wells left to degrade by the companies or individuals that built them. No longer producing usable fossil fuels, these wells nonetheless have a significant impact on the environment, with many of them leaking significant quantities of methane, a powerful greenhouse gas, into the atmosphere. In aggregate, these emissions represent the equivalent of millions of tons of carbon dioxide per year (Williams et al., 2020). In Canada, an estimated 370,000 abandoned wells produce methane equivalent to half a million metric tons of CO2 annually (Williams et al., 2020; ECCC, 2024), while in the U.S., there are an estimated 4 million abandoned wells (Williams et al., 2020), releasing over five million metric tons of CO2 equivalent emissions per year. Despite the small percentage, the high global warming potential of methane underscores the need for improved monitoring and management to reduce environmental impact. Abandoned wells also pose health and safety concerns, in particular by leaching toxic chemicals into the groundwater of surrounding communities (Cahill et al., 2019). It is possible to plug abandoned wells to mitigate the harms associated with them (with so-called “super-emitter” wells an especially high priority (Riddick et al., 2024; Kang et al., 2016)). However, a significant fraction of abandoned wells remain unknown. In Pennsylvania, as much as 90% of abandoned wells are estimated to be unrecorded (Kang et al., 2016). In Canada, abandoned wells have been described as the most uncertain source of methane emissions nationally due to the poor quality of data surrounding them (Williams et al., 2020).
With the advent of large-scale remote sensing datasets and powerful machine learning tools to process them, it has become possible to label and monitor the built environment as never before (Rolf et al., 2024). Many such works have focused on opportunities to use remote sensing to accelerate climate action and environmental protection, and oil and gas infrastructure has increasingly been an object of scrutiny (see e.g. Yang et al. (2013); Sheng et al. (2020)). In this work, we present the first large-scale machine-learning benchmark dataset for pinpointing onshore oil and gas wells, encompassing abandoned, suspended, and active wells. Our main contributions are as follows:
  • We introduce the Alberta Wells Dataset, containing information on over 200k abandoned, suspended, and active onshore oil and gas wells with high-resolution satellite imagery. Our Dataset available at: https://zenodo.org/records/13743323.
@@
NEPU-OWOD V1.0 (Wang et al., 2021) 1,192 432 0.41 Daqing City, China Google Earth
NEPU-OWOD V3.0 (Zhang et al., 2023) 3,749 722 0.48 China & California
Oil Well Dataset (Shi et al., 2021) N/A 5,895 0.26 Daqing City, China
O&G Infrastructure (Guisiano et al., 2024) 630 930 0.15 - 1 Permian Basin, USA
Well Pad Dataset (Ramachandran et al., 2024) 12,490 10,432 0.3-0.7
@@
NEPU-OWS V1.0 (Wu et al., 2023b) N/A 1,200 10 Russia Sentinel-2
NEPU-OWSV2.0 (Wu et al., 2023a) N/A 120 10/20/60 Austin, USA
@@
2 Previous Work
Hundreds of satellites continuously monitor the Earth’s surface, generating petabyte-scale remote sensing datasets (Rolf et al., 2024). With advancements in hardware, the quality of remote sensing images has significantly improved in terms of spatial and temporal resolution. High-quality remote sensing data are available through state-funded projects like Sentinel and Landsat and, more recently, through private enterprises such as Planet Labs (PBC, 2024). Increasingly, machine learning has been used to parse such raw data, including in various applications for tackling climate change (Yang et al., 2013). Benchmark datasets in this area have included tasks in land use and land cover (LULC) estimation (Sumbul et al., 2019), crop classification (Tseng et al., 2021), and forest monitoring (Bountos et al., 2025). Although developed primarily for urban monitoring, the SpaceNet 7 dataset (Van Etten et al., 2021) features a small number of study sites with oil wells.
Within this area of research, an increasing body of work has considered the problem of detecting artifacts associated with oil and gas operations. The detection of oil spills using a combination of remote sensing and machine learning has been widely explored (Wang et al., 2023a; Yang et al., 2022). Recently, the detection of oil and gas infrastructure has also been investigated (Prajapati et al., 2022), with some studies focusing on the goal of estimating methane emissions (Zhu et al., 2022; Omara et al., 2023). The dataset by Sheng et al. (2020) includes 7,066 aerial images, with 149 images of oil refineries. The METER-ML dataset (Zhu et al., 2022) comprises 86,599 georeferenced images in the U.S. labeled for methane sources. The OGIM v1 dataset (Omara et al., 2023) includes 2.6 million point locations of major facilities. A dataset by Chang et al. (2023) features 1,388 images of pipelines in the Arctic, while a dataset by Wang et al. (2023b) includes 3,266 images of heavy-polluting enterprises with 0.25 m resolution.
The problem of detection of oil and gas wells has also been proposed by a number of authors. However, Existing datasets are quite small (500-10,000) and typically are limited to a small region and contain only active wells, limiting their applicability in identifying abandoned or suspended wells as summarized in Table 1. Most of these studies have primarily focused on basic machine learning algorithms for well detection due to the limited sample size.
@@ 3 Alberta Wells Dataset
We introduce the Alberta Wells Dataset for oil and gas well detection. The dataset is drawn from the province of Alberta, Canada, a region with the third-largest oil reserves in the world and a substantial number of oil and gas wells, many of which have been present for over a century. The entire province of Alberta (an area larger than the UK and Germany combined) encompasses a diverse range of geographical zones and is highly diverse for a landlocked region, including prairies, lakes, forests, and mountains. The dataset contains over 188,000 patches of satellite imagery acquired by Planet Labs (PBC, 2024), which we make publicly available, with around 94,000 containing wells covering more than 213,000 individual wells. (More Details in Table 2). Each patch is annotated with labels for both segmentation and bounding box localization. The annotations are based on data from the Alberta Energy Regulator (AER, 2024), quality-controlled by domain experts.
To simulate real-world conditions, we ensure a varied density of wells per image, as highlighted in Figure LABEL:fig:figure1. We also include satellite imagery patches with no wells present from areas nearby to areas with wells, ensuring no overlap between the samples. This balanced dataset maintains an equal distribution of well and non-well images. Table 2 details the total sample count in each dataset split.
3.1 Data Collection, Quality Control & Patch Creation
The Alberta Energy Regulator (AER) oversees the energy industry in the province, ensuring companies adhere to regulations as they develop oil and gas resources. AER publishes AER ST37 (AER, 2024), a monthly list of all wells reported in Alberta, detailing their geographic location, mode of operation, license status, and type of product being extracted, among other attributes. This data provides a metadata (.txt) file and a .shp shape file, where each entry represents a unique geo-location point per site but often contains duplicates. However, this data cannot be used directly because the license status or mode of operation does not always correlate with the actual status of the well and often contains duplicates. Therefore, we work with domain experts to perform quality control as illustrated in Figure 9 (Appendix).
First, we remove duplicate entries from the well metadata, which often contain multiple instances of the same well identified by duplicate license numbers. We resolve these duplicates by retaining the most recent update. A similar approach is applied to the shapefile, where duplicates are resolved using the license date. Afterward, we merge both datasets and filter the data, categorizing the wells as active, abandoned, or suspended based on criteria developed in consultation with domain experts, as shown in Table 3. We check for duplicate location coordinates in the dataset and resolve them by retaining the instance with the latest drill date. Finally, we ensure all the well instances in the dataset are within Alberta’s boundaries. The raw metadata file has around 637,000 instances, while the surface hole geometry file has around 512,000 instances. After filtering and performing quality control on the datasets with domain experts, we have around 217,000 instances. We then calculate the geographical bounds covered by the well instances across the province and divide the region into non-overlapping square image patches, each covering an area of 1.1025 sq km (with sides of 1050m). These images include various numbers of individual wells (see Figure  LABEL:fig:figure1), and we ensure that an approximately equal number of patches exist with and without wells. As a result of this process, some samples were excluded due to being located outside Alberta’s geographical boundaries, leading to a final total of approximately 213,000 well instances in the dataset patches.
@@ 3.2 Dataset Splitting
To create a well-distributed dataset that represents various geographical regions and offers a diverse dataset for evaluation, we developed a splitting algorithm (Algorithm 1) which focuses on balancing regions, not individual examples, ensuring that both the training and test sets reflect a diverse range of regions from Alberta’s varied landscape. This approach preserves dataset diversity and simulates real-world conditions where imbalances are common.
@@
This method involves forming small clusters $k_{1i}$ of nearby well patches based on their centroids as illustrated in Figure LABEL:fig:figuresplit (a). These small clusters are then grouped into larger, non-intersecting super-clusters $k_{2i}$, with each super-cluster representing a city or larger geographical area. The formation of super-clusters involves calculating a centroid for each $k_{1i}$ cluster based on the centroids of the well patches it contains as illustrated in Figure LABEL:fig:figuresplit (b). By clustering wells in this manner, we ensure that $k_{1i}$ clusters group wells from nearby localities together, while $k_{2i}$ clusters group wells from the same geographic region as illustrated in Figure LABEL:fig:figuresplit (c). Thus, each $k_{2i}$ cluster represents a geographic distribution, with each $k_{1i}$ cluster within it representing a sample of that distribution. To ensure a diverse and well-distributed evaluation and testing of our machine learning model, we select the $k_{1i}$ clusters with the two fewest well instances from each $k_{2i}$ super-cluster for inclusion in the evaluation and test sets. This approach ensures a diverse representation of the dataset as observed in Figure LABEL:fig:figuresplit (d). Moreover, we maintain an equal distribution of well and non-well patches. In cases of imbalance in non-well images, we exclude such patches from the contributing $k_{1i}$ clusters as specified in Algorithm 1. For imbalances in well images, we sample non-well patches that are not part of any other clusters.
@@ 3.3 Satellite Imagery Acquisition & Label Creation
We used PlanetScope-4-Band imagery (PBC, 2024) featuring RGB and Near Infrared bands to represent satellite images of the region with a medium resolution of about 3 meters per pixel. PlanetScope, a product of Planet Labs, consists of approximately 130 satellites that can image the entire Earth’s land surface daily, collecting up to 200 million sq. km of data each day. We obtained Surface Reflectance imagery, which is offset-corrected, flat-field-corrected, ortho-rectified, visually processed, and radio-metrically corrected. These processes ensure consistency across varying atmospheric conditions and minimize uncertainty in spectral response over time and location, making the data ideal for various applications.
We choose Planet Labs data over other alternatives since it is updated daily, making it possible to pick a consistent time for all the images, which is important for training dataset consistency. It also provides multispectral imagery (4-band: RGB+Near Infrared), and the Near Infrared band is a useful addition (as we later verify in Table 7) since certain features, like ground depressions indicating well sites, may be more detectable in this band. Lastly, while other alternatives may be limited in remote regions, Planet’s global satellite constellation ensures more consistent coverage. All the imagery we use is made publicly available in the dataset. To ensure the highest quality, we selected images with no cloud cover. The images were acquired by Planet satellites within a timeframe that aligns with the well-location data from AER. We obtained satellite images for each sample based on geographical coordinates, ensuring an intersection between the actual area of interest and the acquired imagery.
@@
We frame the task of identifying wells as both an object detection and segmentation task. In a real-world setting, high performance in either detection or segmentation would be sufficient for detection of wells; therefore we test both framings since it is not a priori clear which may be best suited to remote sensing algorithms. For each image patch, as shown in Figure 7, we generated corresponding segmentation maps and object detection annotations for all known wells in the image based on the point labels provided in the AER data. For binary segmentation, we annotated each well site with a circle to match the teardrop shape typical for well sites. We standardized the diameter of a well site to a value of 90 meters (such sites typically range from 70 to 120 meters in diameter). We used the same scale to define bounding boxes in the object detection task, following the COCO (Lin et al., 2014) format for annotations. The overlap in ground truth bounding boxes for some of the wells in Figures 7,8,17 and 18 reflects the clustering of multiple wells in densely developed oil and gas sites, where the spatial overlap of wells and infrastructure is common. (Note that this is a characteristic of the data, not a limitation of our quality control strategy.) Additionally, we created multi-class segmentation maps, where each class represents a different state of the well (active, suspended, or abandoned), and included this information in the object detection annotations. (We do not perform multi-class segmentation experiments here, but it is possible that future researchers may find this task useful.)
Figure 7: Sample image patches from our dataset includes examples with no wells, two wells, and multiple wells. Additionally, we present qualitative results with predictions generated by our Segmentation U-Net (ResNet50) and DETR ResNet50 model.